Consider the test of H0:σ^2=7 against H1:σ^2>7. What is the critical value for the test statistic X02 for the significance level α=0.05 and sample size n=19?

Give your answer with two decimal places (e.g. 98.76).

Enter your answer in accordance to the question statement

Thank you!
Answer #1

We have given,

Sample size =n=19

Degree of freedom =n-1=19-1=18

Chi square critical value for right tailed test =28.87...........by using chi square critical value table or Excel =CHIINV(0.05,18)